# HALT: Hallucination Assessment via Latent Testing

### v1

@@ Abstract
Hallucination in large language models (LLMs) can be understood as a failure of faithful readout: although internal representations may encode uncertainty about a query, decoding pressures still yield a fluent answer. We propose lightweight residual probes that read hallucination risk directly from intermediate hidden states of question tokens, motivated by the hypothesis that these layers retain epistemic signals that are attenuated in the final decoding stage. The probe is a small auxiliary network whose computation is orders of magnitude cheaper than token generation and can be evaluated fully in parallel with inference, enabling near-instantaneous hallucination risk estimation with effectively zero added latency in low-risk cases. We deploy the probe as an agentic critic for fast selective generation and routing, allowing LLMs to immediately answer confident queries while delegating uncertain ones to stronger verification pipelines. Across four QA benchmarks and multiple LLM families, the method achieves strong AUROC and AURAC, generalizes under dataset shift, and reveals interpretable structure in intermediate representations, positioning fast internal uncertainty readout as a principled foundation for reliable agentic AI.
@@ 1 Introduction
We propose a lightweight hallucination-detection method based on question intermediate representations that predicts hallucination risk efficiently and accurately. The additional computation introduced by our method is less than 1‰ of that required to generate a single token, so our approach adds negligible cost to generation. Moreover, the detector can be evaluated in parallel with inference, introducing no extra latency to the user. In essence, we aim to build a “lie detector” or “mind reader” for LLMs.
With an effective hallucination detector, we can build an LLM router that adaptively allocates computation. Given a user query, the default LLM begins generating a response while our detector estimates the hallucination risk in parallel. Because the detector only requires the question’s intermediate representations, it can run concurrently with generation. If the predicted risk is low, the system returns the generated response as usual; otherwise, it routes the query to a slower but more reliable pipeline (e.g., a stronger model, reasoning-augmented generation, cross-model verification, or retrieval-augmented generation (RAG)). In the low-risk case, this design introduces zero additional latency; in the fallback case, the extra delay is less than the time to generate a single token.
Figure 1: Comparison of hallucination handling strategies. Traditional detection pipelines wait for generation to complete before checking for hallucinations, doubling latency in the fallback case. Our method evaluates hallucination risk in parallel with generation, enabling zero-latency responses for confident queries while routing uncertain ones to stronger models.
@@
Based on the above observations, we hypothesize that intermediate representations encode uncertainty signals. This is analogous to how students taking an exam can sense “I don’t know this” for a question, yet would never write that on the answer sheet. LLMs exhibit a similar pattern because they are trained to provide an answer whenever possible. This motivates us to build a small detector that captures this internal sense of confusion. Moreover, this analogy suggests that intermediate layers should be more informative than the final output, just as the internal feeling differs from what appears on the answer sheet.
In our experiments, we indeed find that using representations from intermediate layers is often more effective than using the final-layer representations. We attribute this to the fact that the final layer has already been decoded into the token space, leading to significant information loss. In other words, the model contains internal features related to confidence or hallucination, but these features are unnecessary for output generation, so they may be discarded in the last few layers.
Furthermore, we only use the representations aligned to the question as input to the detector. This choice is driven by the need to balance detection quality and response latency. In our experiments, we find that question-only representations are already sufficient for accurate detection. Including answer representations yields a marginal improvement but incurs substantial latency overhead.
Our main contributions are as follows.
  • We propose a lightweight hallucination detection method that incurs negligible computational overhead and can be evaluated in parallel with inference.
  • Building on this detector, we develop a pipeline that routes queries with high hallucination risk to stronger models to improve generation quality without increasing response latency or token generation cost. Preliminary experiments show that this pipeline improves answer correctness by xxx.
  • We conduct extensive ablations and identify several properties of intermediate representations. (a) Intermediate-layer representations are more effective than the final layer for hallucination detection. (b) While question-aligned representations are often sufficient for accurate risk estimation, including answer tokens improves our method.
2 Related Work
@@
Ours No Need Middle Hidden States Question
@@ 3 Methodology
We formulate hallucination detection as a knowledge prediction task. Given a user query $q$, our goal is to estimate whether the LLM contains the correct answer before or during generation. During the forward process, the LLM produces hidden representations $h_{l}\in\mathbb{R}^{N\times D}$ for each layer $l\in[0,L]$, where $N$ is the sequence length, $D$ is the hidden dimension, and $L$ is the number of transformer layers. We use a lightweight auxiliary network $g_{\theta}$ that takes the hidden representation $h$ as input and outputs a scalar score
@@
The detector relies on the LLM’s hidden representations to infer whether the model is confident about a given query. In principle, representations from any transformer layer can be used as input. We choose to extract features from intermediate layers rather than the final one. As discussed earlier, the last layer is optimized for next-token prediction and may discard signals unrelated to generation. Intermediate layers, by contrast, retain richer representations that capture the model’s internal uncertainty.
We also restrict the detector’s input to hidden states corresponding only to the question tokens. Formally, the feature we use can be represented as
@@
where $h_{l}$ denotes the hidden representation of the query(question) $q$ at layer $l$. This choice reduces computation and avoids the need to process the model’s generated text. More importantly, it allows the hallucination likelihood to be estimated before any answer is produced. Methods that depend on analyzing the generated output must first wait for the model to finish responding, then, if the answer is found unreliable, invoke a stronger model to regenerate it, effectively doubling the latency. Our approach avoids this inefficiency by evaluating the question representations during the original forward pass, so the system can proactively switch to a more reliable model and introduce almost no additional delay. For factual benchmarks, we achieve performance comparable to or better than output-based methods, using only the question representations without waiting for the model’s response.
@@
We construct the training data from question-answering benchmarks that include reference answers. For each question $q$ with the standard answer $a^{\ast}$, the target LLM generates one answer $a$. We then obtain a correctness label by using an external judge to compare $(q,a)$ with $a^{\ast}$. In particular, we use gpt4o for to produce more accurate labels than rule-based evaluation methods. In principle, multiple answers can be sampled for each question to estimate the probability of correctness. Still, in practice, we generate only one answer per question due to time and cost limitations. The resulting label is $y\in\{0,1\}$, where $y=1$ means the generated answer is correct and $y=0$ means it is hallucinated or incorrect.
Given the detector output $p\in[0,1]$, we employ a binary cross-entropy objective. For a dataset with $M$ examples $\{(q_{i},y_{i})\}_{i=1}^{M}$ and detector scores $p_{i}$, we minimize
$\mathcal{L}=-\frac{1}{M}\sum_{i=1}^{M}\big(y_{i}\log p_{i}+(1-y_{i})\log(1-p_{i})\big).$
In closed-book question answering, the model generates answers solely from its parametric knowledge without access to external sources. An incorrect answer in this setting indicates that the model has produced content inconsistent with factual ground truth, which aligns with the definition of factual hallucination. We therefore treat correctness labels as a proxy for hallucination detection, following prior work in this area (Du et al., 2024; Farquhar et al., 2024).
Network Architecture (Hallucination Detector)
@@
For the MLP architecture, the input must have a fixed length. Therefore the hidden representation $h_{l}\in\mathbb{R}^{N\times D}$ is first compressed into a single vector in $\mathbb{R}^{D}$. We consider several simple aggregation methods, including mean pooling, max pooling, and using the hidden state of the last token, which often corresponds to the question mark. We also consider applying principal component analysis (PCA) to the hidden representations and retain the top $n$ principal components. The resulting pooled vector or concatenated principal components are then fed into an MLP that predict the confidence score $p$.
@@ 4 Experiments
Table 3: Summary of four QA benchmarks used for hallucination detection evaluation.
@@
LLaMA 2 Chat 7B HaloScope 77.40 67.0 79.6 72.6
@@
Qwen-2.5-7B-Instruct HaloScope 85.5 70.3 81.1 80.4
@@
Gemma-3-4b-it HaloScope 76.5 79.62 77.7 76.7
@@
Our experiments use four open-domain question answering benchmarks, namely TriviaQA (Joshi et al., 2017), NQ-Open (Kwiatkowski et al., 2019), MMLU-Pro (Wang et al., 2024), and WebQuestions (Berant et al., 2013), summarized in Table 3. TriviaQA contains general knowledge trivia questions in an unfiltered no-context setting. NQ-Open consists of naturally occurring queries from Google Search paired with Wikipedia-based answers. MMLU-Pro provides multiple-choice questions spanning 14 professional and academic subjects, often requiring multi-step reasoning. WebQuestions includes entity-centric questions originally collected from the Google Suggest API. We evaluate our approach on three model families, namely LLaMA-2 (Touvron et al., 2023), Qwen2.5 (Yang et al., 2024), and Gemma-3 (Team, 2025). Ablation studies on model scale are conducted using the 7B, 14B, and 32B instruct variants of Qwen2.5. Detailed statistics on correctness labels for each model-dataset pair are provided in Table 5.
@@
In addition, accuracy and AURAC (area under the rejection–accuracy curve) are also reported. The RAC plots accuracy against coverage, where samples are ranked by confidence and coverage denotes the fraction of samples retained. AURAC evaluates whether the model’s confidence is reliable, with higher values indicating a stronger alignment between correctness and confidence. Figure 3 provides a concrete example of a RAC curve, which may aid understanding of this metric.
@@
We use GPT-4o to obtain correctness labels for LLM-generated answers. For each question, we prompt GPT-4o with the original question, the ground-truth answer, and the LLM-generated answer, asking it to return a binary correctness label (see Appendix A for the full prompt). This approach provides more robust evaluation than rule-based string matching, particularly for open-ended questions where semantically equivalent answers may differ in surface form.
@@ 4.1 Main Results
Table 4 presents the main results. Our method achieves the best AUROC on 10 out of 12 model-dataset combinations. Qwen-2.5-7B-Instruct shows the strongest performance, where our approach outperforms all baselines across all four benchmarks, with gains of up to 8.5 points on TriviaQA and 8.2 points on NQ-Open. On LLaMA 2 Chat 7B and Gemma-3-4b-it, we observe similar improvements, surpassing prior methods by 7.7 and 12.3 points on TriviaQA respectively. Using answer token representations (Ours answer) consistently outperforms question token representations (Ours question) across all settings, suggesting that answer tokens encode richer information about the model’s factual confidence. The two exceptions occur on MMLU-Pro for LLaMA 2 and Gemma-3, where HaloScope achieves slightly higher AUROC. We attribute this to two factors: (1) a format mismatch, as TriviaQA, NQ-Open, and WebQuestions are open-ended QA tasks while MMLU-Pro is a multiple-choice benchmark where the model selects from predefined options rather than generating free-form answers; and (2) label imbalance, as LLaMA 2 and Gemma-3 achieve only 15.4% and 27.1% accuracy on MMLU-Pro respectively (see Table 5), resulting in limited positive samples for training the detector. Overall, these results demonstrate that intermediate-layer representations provide a strong signal for hallucination detection, competitive with or superior to methods that require sampling multiple outputs.
@@
We evaluate whether our detector generalizes under distribution shift. Unless otherwise specified, we extract hidden states from layer 19 of Qwen-2.5-7B-Instruct. We train the detector on each of the four datasets and evaluate it on the remaining datasets. We also train on the union of all datasets (All) and test on each dataset. Figure 2 reports the AUROC for each train–test pair.
The detector exhibits strong out-of-distribution generalization. Most off-diagonal entries remain in the 75–93 AUROC range, indicating that the intermediate representations capture dataset-agnostic signals predictive of hallucination rather than dataset-specific patterns. Training on the union (All) yields a robust model that achieves AUROC of 94.20 on TriviaQA, 88.78 on NQ-Open, 86.78 on MMLU-Pro, and 87.17 on WebQuestions. While All does not uniformly dominate the best single-source model for every target, it provides competitive and stable performance across all datasets, reducing sensitivity to the particular source distribution. These results support our conclusion that the learned features are largely transferable across different QA domains.
@@ 4.2 Accuracy after Hallucination Removed
Figure 3: Rejection-Accuracy Curve (RAC) for Qwen-2.5-7B on TriviaQA using only question token representations. As we reject samples with low confidence, accuracy increases from 85.7% at full coverage to 97.7% at 40% coverage. AURAC = 0.8753 indicates strong alignment between confidence and correctness.
Beyond binary hallucination detection, our method can be used for selective prediction: abstaining from answering when the model is likely to hallucinate. Crucially, because we use only question token representations, the detector can evaluate hallucination risk before the model generates any answer, introducing zero additional latency. We use the detector’s output probability as a confidence score and reject samples below a threshold. Figure 3 shows the rejection-accuracy curve (RAC) for Qwen-2.5-7B on TriviaQA, where we vary the confidence threshold and measure accuracy on the retained samples.
At full coverage, the model achieves 85.7% accuracy. By rejecting the 20% least confident samples, accuracy rises to 91.6%; at 60% coverage, it reaches 95.6%; and at 40% coverage, it climbs to 97.7%. The area under the RAC (AURAC = 0.8753) quantifies this strong alignment between detector confidence and correctness.

### v2

# DRIFT: Detecting Representational Inconsistencies for Factual Truthfulness

@@ Abstract
LLMs often produce fluent but incorrect answers, yet detecting such hallucinations typically requires multiple sampling passes or post-hoc verification, adding significant latency and cost. We hypothesize that intermediate layers encode confidence signals that are lost in the final output layer, and propose a lightweight probe to read these signals directly from hidden states. The probe adds less than 0.1% computational overhead and can run fully in parallel with generation, enabling hallucination detection before the answer is produced. Building on this, we develop an LLM router that answers confident queries immediately while delegating uncertain ones to stronger models. Despite its simplicity, our method achieves SOTA AUROC on 10 out of 12 settings across four QA benchmarks and three LLM families, with gains of up to 13 points over prior methods, and generalizes across dataset shifts without retraining.
@@ 1 Introduction
Existing methods fall into two categories. Sampling-based approaches, such as Semantic Entropy (Farquhar et al., 2024), generate multiple answers and compute semantic uncertainty, but require multiple forward passes, incurring a high computational cost that is incompatible with real-time requirements. Representation-based approaches, such as HaloScope (Du et al., 2024), train classifiers on final-layer hidden states to predict hallucination risk, but final-layer representations have already been projected toward the token space, potentially discarding internal confidence signals.
We propose a lightweight hallucination-detection method based on intermediate representations that predicts hallucination risk efficiently and accurately. Unlike prior representation-based methods that assume a fixed low-dimensional truth subspace, our approach detects representational instability before projection to token space. The additional computation introduced by our method is less than 0.1% of that required to generate a single token, so our approach adds negligible cost to generation. Moreover, the detector can be evaluated in parallel with inference, without adding extra latency for the user. In essence, we aim to build a “lie detector” or “mind reader” for LLMs.
With an effective hallucination detector, we can build an LLM router that adaptively allocates computation. Given a user query, the default LLM begins generating a response while our detector estimates the risk of hallucination in parallel. Because the detector can rely solely on the question’s intermediate representations, it can run concurrently with generation. If the predicted risk is low, the system returns the generated response as usual; otherwise, it routes the query to a slower but more reliable pipeline (e.g., a stronger model, reasoning-augmented generation, cross-model verification, or retrieval-augmented generation (RAG)). In the low-risk case, this design introduces zero additional latency; in the fallback case, the extra delay is less than the time to generate a single token.
Figure 1: Comparison of hallucination handling strategies. Traditional detection pipelines wait for the generation to complete before checking for hallucinations, doubling latency in the fallback case. Our method evaluates hallucination risk in parallel with generation, enabling zero-latency responses for confident queries and supporting selective prediction and compute allocation while routing uncertain ones to stronger verification pipelines.
@@
Based on the above observations, we hypothesize that intermediate representations encode uncertainty signals. This is analogous to how students taking an exam can sense, for a question, “I don’t know this” yet would never write that on the answer sheet. LLMs exhibit a similar pattern because they are trained to provide an answer whenever possible. This motivates us to build a small detector that captures this internal sense of confusion. Moreover, this analogy suggests that intermediate layers should be more informative than the final output, just as the internal feeling differs from what appears on the answer sheet.
In our experiments, we find that using representations from intermediate layers is often more effective than using the final-layer representations. As noted earlier, final-layer representations lose information during projection to token space. This occurs because the model may contain internal features related to confidence, but since these features are unnecessary for next-token prediction, they are discarded in the last few layers.
Moreover, recent work shows that larger models perform richer latent computation, and their chain-of-thought outputs can be less faithful to internal reasoning (Chen et al., 2025). This suggests that the gap between intermediate representations and surface outputs may widen with model scale. We therefore hypothesize that our method should be more effective when applied to larger models.
Furthermore, our method supports two operating modes with different latency-accuracy trade-offs. Using only question representations enables prediction before the answer is generated, allowing proactive model switching with minimal latency overhead. While answer representations yield higher accuracy (as shown in Table 4), question-only detection already outperforms prior baselines in most settings, making it a practical choice when low latency is prioritized.
@@
Ours No Need Middle Hidden States Question/Answer
Our main contributions are as follows.
  • We propose a lightweight hallucination detector that reads uncertainty signals from intermediate hidden states. The detector incurs less than 0.1% computational overhead and can be evaluated in parallel with inference, introducing zero latency.
  • On four QA benchmarks across three LLM families, our method achieves SOTA AUROC on 10 out of 12 settings, with gains of up to 13 points over prior methods, and exhibits strong out-of-distribution generalization across datasets.
  • We develop an LLM router that leverages question-only predictions to route high-risk queries to stronger models before answer generation, improving answer correctness without increasing response latency.
  • Through extensive ablations, we identify that (a) intermediate-layer representations outperform final-layer representations, and (b) our method benefits more from larger models, consistent with our hypothesis.
2 Related Work
@@ 3 Methodology
We formulate hallucination detection as a knowledge prediction task. Given a user query $q$, our goal is to estimate whether the LLM can correctly answer it before or during generation.
We focus on closed-book question answering, where the model generates answers solely from its parametric knowledge without access to external sources. An incorrect answer in this setting indicates that the model has produced content inconsistent with factual ground truth, which aligns with the definition of factual hallucination. As in prior work (Du et al., 2024; Farquhar et al., 2024), we treat correctness labels as a proxy for hallucination detection.
During the forward process, the LLM produces hidden representations $h_{l}\in\mathbb{R}^{N\times D}$ for each layer $l\in[0,L]$, where $N$ is the sequence length, $D$ is the hidden dimension, and $L$ is the number of transformer layers. We use a lightweight auxiliary network $g_{\theta}$ that takes the hidden representation $h$ as input and outputs a scalar score
@@
The detector takes hidden representations from a single intermediate layer as input. We select the layer index as a hyperparameter and find that intermediate layers outperform the final layer (see Figure 4). Formally, the feature we use can be represented as
@@
where $h_{l}$ denotes the hidden representation of the query $q$ at layer $l$.
By default, we use only the hidden states corresponding to question tokens, which allows the detector to run before answer generation begins. When answer tokens are available, including them improves accuracy at the cost of increased latency.
@@
We construct the training data from question-answering benchmarks that include reference answers. For each question $q$ with the standard answer $a^{\ast}$, the target LLM generates one answer $a$. We then obtain a correctness label by using an external judge to compare $(q,a)$ with $a^{\ast}$. In particular, we use GPT-4o to produce more accurate labels than rule-based evaluation methods. In principle, multiple answers can be sampled for each question to estimate the probability of correctness. Still, in practice, we generate only one answer per question due to time and cost limitations. The resulting label is $y\in\{0,1\}$, where $y=1$ means the generated answer is correct and $y=0$ means it is hallucinated or incorrect. Given the detector output $p\in[0,1]$, we train the network by minimizing the binary cross-entropy loss between $p$ and the correctness label $y$.
Detector Architecture
@@
For the MLP architecture, the input must have a fixed length. Therefore the hidden representation $h_{l}\in\mathbb{R}^{N\times D}$ is first compressed into a single vector in $\mathbb{R}^{D}$. We consider several simple aggregation methods, including mean pooling, max pooling, and using the hidden state of the last token, which often corresponds to the question mark. We also consider applying principal component analysis (PCA) to the hidden representations and retaining the top $n$ principal components. The resulting pooled vector or concatenated principal components are then fed into an MLP that predicts the confidence score $p$.
@@ 4 Experiments
Table 3: Summary of four QA benchmarks used for hallucination detection evaluation. MCQ = multiple-choice question.
@@
LLaMA 2 Chat 7B HaloScope(question) 66.92 66.97 79.56 72.62
HaloScope(answer) 77.40 67.05 74.83 77.03
@@
Qwen-2.5-7B-Instruct HaloScope(question) 85.48 70.33 81.08 80.43
HaloScope(answer) 86.54 85.84 77.56 78.56
@@
Gemma-3-4b-it HaloScope(question) 76.52 79.62 77.73 76.70
HaloScope(answer) 70.13 85.84 77.56 75.72
@@
Our experiments use four open-domain question answering benchmarks, namely TriviaQA (Joshi et al., 2017), NQ-Open (Kwiatkowski et al., 2019), MMLU-Pro (Wang et al., 2024), and WebQuestions (Berant et al., 2013), summarized in Table 3. TriviaQA contains general knowledge trivia questions in an unfiltered, no-context setting. NQ-Open consists of naturally occurring queries from Google Search paired with Wikipedia-based answers. MMLU-Pro provides multiple-choice questions spanning 14 professional and academic subjects, often requiring multi-step reasoning. WebQuestions includes entity-centric questions originally collected from the Google Suggest API. We evaluate our approach on three model families, namely LLaMA-2 (Touvron et al., 2023), Qwen2.5 (Yang et al., 2024), and Gemma-3 (Team, 2025). Ablation studies on model scale are conducted using the 7B, 14B, and 32B instruct variants of Qwen2.5. Detailed statistics on correctness labels for each model-dataset pair are provided in Table 5.
@@
In addition, accuracy and AURAC (area under the rejection–accuracy curve) are also reported. The RAC plots accuracy against coverage, where samples are ranked by confidence, and coverage denotes the fraction of samples retained. AURAC assesses whether the model’s confidence is reliable, with higher values indicating greater alignment between correctness and confidence. Figure 3 provides a concrete example of an RAC curve, which may aid understanding of this metric.
@@
Rule-based correctness labeling methods (e.g., string matching) are unreliable for two reasons. First, for open-ended questions, semantically equivalent answers may differ in surface form. Second, when models produce chain-of-thought outputs, intermediate reasoning steps may mention the correct answer, which can lead to mislabeling. We therefore adopt a more expensive but reliable approach, using GPT-4o to obtain correctness labels. For each question, we prompt GPT-4o with the original question, the ground-truth answer, and the LLM-generated answer, asking it to return a binary correctness label (see Appendix A for the full prompt). This approach provides a more robust evaluation than rule-based matching.
@@ 4.1 Main Results
Table 4 presents the main results. For each method-model-dataset-feature combination, we use the best-performing layer to ensure a fair comparison. Our method achieves the best AUROC on 10 out of 12 model-dataset combinations. Qwen-2.5-7B-Instruct shows the strongest performance, with our approach outperforming all baselines across all four benchmarks, achieving gains of up to 13.0 points (question-only on NQ-Open) and 9.5 points (question-plus-answer on MMLU-Pro). On LLaMA 2 Chat 7B and Gemma-3-4b-it, we observe similar improvements, surpassing prior methods by 7.7 and 12.3 points on TriviaQA, respectively. The two exceptions occur on MMLU-Pro for LLaMA 2 and Gemma-3, where HaloScope achieves slightly higher AUROC. We attribute this to two factors. First, a format mismatch, as TriviaQA, NQ-Open, and WebQuestions are open-ended QA tasks while MMLU-Pro is a multiple-choice benchmark where the model selects from predefined options rather than generating free-form answers. Second, label imbalance: for example, LLaMA 2 and Gemma-3 achieve only 15.4% and 27.1% accuracy on MMLU-Pro, respectively (see Table 5), resulting in limited positive samples for training the detector. Using answer token representations (Ours answer) consistently outperforms question token representations (Ours question) across all settings, suggesting that answer tokens encode richer information about the model’s factual confidence. Overall, these results demonstrate that intermediate-layer representations provide a strong signal for hallucination detection, competitive with or superior to methods that require sampling multiple outputs.
@@
4.2 Out-of-Distribution Generalization
In practice, the distribution of user queries may differ from the training data. We therefore evaluate whether our detector generalizes under distribution shift. We train the detector on each of the four datasets and evaluate it on the remaining datasets. We also train on the union of all datasets (All) and test on each dataset. Unlike the main experiments, where we select the best layer per setting, here we fix layer 19 of Qwen-2.5-7B-Instruct for all experiments, since tuning the layer on the target dataset would violate the out-of-distribution setting. Figure 2 reports the AUROC for each train-test pair.
As the heatmaps show, the detector exhibits strong out-of-distribution generalization in both settings. For question-only detection, most off-diagonal entries remain in the 72–88 AUROC range; for question-plus-answer, this improves to 75–93. Notably, training on the union of all datasets (All) yields strong and stable performance across all four targets, indicating that the learned features are dataset-agnostic rather than overfitting to source-specific patterns. These results suggest that practitioners can train a single detector on available labeled data and deploy it across diverse query distributions without retraining.
@@ 4.3 Accuracy after Hallucination Removed
Figure 3: Rejection-Accuracy Curve (RAC) for Qwen-2.5-7B on TriviaQA using question token representations. Rejecting low-confidence samples substantially improves accuracy on retained samples.
Beyond binary hallucination detection, our method can be used for selective prediction: abstaining from answering when the model is likely to hallucinate. In this experiment, we use only question token representations, allowing the detector to evaluate hallucination risk before the model generates any answer, introducing zero additional latency. We use the detector’s output probability as a confidence score and reject samples below a threshold. Figure 3 shows the rejection-accuracy curve (RAC) for Qwen-2.5-7B on TriviaQA, where we vary the confidence threshold and measure accuracy on the retained samples.
At full coverage, the model achieves 85.7% accuracy. By rejecting the 20% least confident samples, accuracy rises to 91.6%. At 60% coverage, accuracy reaches 95.6%, and at 40% coverage, it climbs to 97.7%. The area under the RAC (AURAC = 0.8753) quantifies this strong alignment between detector confidence and correctness.